I got a question for you
State the maximum value of the function y= 3x^2 - 18x +4 (by finding the y coordinate of its vertex)

Answer by lwsshak3(11628)

State the maximum value of the function:
y= 3x^2 - 18x +4
complete the square:
y=3(x^2-6x+9)-27+4
y=3(x-3)^2-23
This is a parabola that opens upward. (Has a minimum)
Its standard form of equation: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of vertex.
Function has a minimum of -23 at x=3
Range: (-23,∞)